OpenStudy (anonymous):

Use polynomial long division to divide 4x^3 + x^2 + 17x – 15 by -x^2 – x – 5

OpenStudy (anonymous):

So you want to find the polynomial which, when multiplied by (-x^2-x-5), equals the big one. Take the highest power -x^2. What will we multiply this by to get 4x^3? Yes, -4x. So (-x^2-x-5)*(-4x) = 4x^3+4x^2+20x. The first term is correct, but the second is 4x^2 when we only want x^2, so we need to add -3x^2 to what we already have. So AGAIN we look at the largest term -x^2. What will we multiply this by to get -3x^2? Yes, 3. So (-x^2-x-5)*(-4x) = 4x^3+4x^2+20x. So (-x^2-x-5)*(3) = -3x^2-3x -15. So (-x^2-x-5)*(-4x+3) =4x^3 + x^2+17x -15 Continue this process until you are finished, or you are left with a remainder.

OpenStudy (anonymous):

thank you!!